I know that I need three equations to solve the system, but it seems like I don't have enough information to solve. I can do x+y+z=40 for one equation. But I'm not sure how to get the other two equations. Thanks for the help.
The college bookstore sells 3 types of calculators. They have a total of 40 calculators in stock. The total amount of revenue from these calculators will be $3590. The most expensive type costs $75 more than the cheapest. Find the cost of each type.

Answer by venugopalramana(3286)

I know that I need three equations to solve the system, but it seems like I don't have enough information to solve. I can do x+y+z=40 for one equation. But I'm not sure how to get the other two equations. Thanks for the help.
The college bookstore sells 3 types of calculators. They have a total of 40 calculators in stock. The total amount of revenue from these calculators will be $3590. The most expensive type costs $75 more than the cheapest. Find the cost of each type.
THERE ARE NOT 3 BUT 4 UNKNOWNS HERE INCLUDING THE RATES
LET X =NUMBER OF CHEAPEST CALCULATORS SOLD AT RATE OF A PER CALCULATOR.
LET Y =NUMBER OF COSTLIEST CALCULATORS SOLD AT RATE OF A+75 PER CALCULATOR.
LET 40-X-Y =NUMBER OF MIDCOST CALCULATORS SOLD AT RATE OF B PER CALCULATOR.
SO THE EQUATIONS IS
AX+(A+75)Y+B(40-X-Y)=3590
HENCE THEWRE ARE 4 UNKNOWNS A,X,Y,B WITH 1 EQN.BUT THERE IS ONE CONDITION THAT X AND Y SHOULD BE INTEGERS AS THEY ARE NUMBER OF CALCULATORS SOLD.
WE CAN ONLY TRY BY GUESS OR USE ADVANCED NUMBER THEORY TECHNIQUES TO REDUCE NUMBER OF TRIALS TO ANSWER.
I DO NOT KNOW YOUR COURSE OF STUDY OR YOUR BACK GROUND KNOWLEDGE TO DWELL ON THESE THINGS.PLEASE INFORM .OR...
GET BACK TO YOUR TEACHER WITH THIS REPORT.